Hi
the sides of a rectangular park are in the ratio of 3:2and its area is 3750m^2.what is the cost of
fencing it at $50 a metre.
thank u

Answer by macston(5194)

.
L=length; W=width; A=area; P=perimeter
.
2L=3W
L=(3/2)W=1.5W
.
A=LW
3750m^2=(1.5W)(W)
3750m^2=1.5W^2
2500m^2=W^2
50m=W
Width of the park is 50 meters.
.
L=1.5W=1.5(50m)=75m
Length of the park is 75 meters.
.
Amount of fence needed = P
P=2(L+W)
P=2(75m+50m)
P=2(125m)
P=250m
It will take 250 meters to fence the park
.
Total cost=P($50)=(250m)($50/m)=$12500
.
ANSWER: Total cost of fencing the park would be $12500.
.